OpenStudy (yanasidlinskiy):

Solving equations and Formulas for a variable. x+5y = 45 Solve for y.

OpenStudy (yanasidlinskiy):

My answer is \[y =- \frac{ 1 }{ 5 } x + 9\]

OpenStudy (misty1212):

looks good to me!

OpenStudy (yanasidlinskiy):

How is that wrong?!

OpenStudy (bloomlocke367):

You have to isolate y. To do that, first subtract x from both sides. \(5y=45-x\)

OpenStudy (yanasidlinskiy):

Well, I got it wrong so I do not understand on what I'm missing.

OpenStudy (misty1212):

what else could it be?

OpenStudy (bloomlocke367):

Then, divide both sides by 5.

OpenStudy (misty1212):

\[x+5y=45\\ 5y=-x+45\\ y=-\frac{x}{5}+9\]